The solve the system using the Linear combination mathod.
21x-8y=-1
9x+5y=8

Answer by kikerkins(17)

21x-8y=-1
9x+5y=8

Let's solve for X first
5(21x-8y=-1)
8(9x+5y=8)
Distribute.
105X-40Y=-5
72X+40Y=64
Add down
177X=59 ( the Y's cancel)

X=59/177
Now plug 59/177 in for X
9(59/177)+5y=8
513/177+5Y=8
5Y=301/59
Y=301/295